# SI-MIL: Taming Deep MIL for Self-Interpretability in Gigapixel Histopathology

### v1

@@ Abstract
Introducing interpretability and reasoning into Multiple Instance Learning (MIL) methods for Whole Slide Image (WSI) analysis is challenging, given the complexity of gigapixel slides. Traditionally, MIL interpretability is limited to identifying salient regions deemed pertinent for downstream tasks, offering little insight to the end-user (pathologist) regarding the rationale behind these selections. To address this, we propose Self-Interpretable MIL (SI-MIL), a method intrinsically designed for interpretability from the very outset. SI-MIL employs a deep MIL framework to guide an interpretable branch grounded on handcrafted pathological features, facilitating linear predictions. Beyond identifying salient regions, SI-MIL uniquely provides feature-level interpretations rooted in pathological insights for WSIs. Notably, SI-MIL, with its linear prediction constraints, challenges the prevalent myth of an inevitable trade-off between model interpretability and performance, demonstrating competitive results compared to state-of-the-art methods on WSI-level prediction tasks across three cancer types. In addition, we thoroughly benchmark the local- and global-interpretability of SI-MIL in terms of statistical analysis, a domain expert study, and desiderata of interpretability, namely, user-friendliness and faithfulness.
Code: https://github.com/bmi-imaginelab/SI-MIL/tree/main
@@ 1 Introduction
Figure 1: Unlike conventional MIL, SI-MIL co-learns from deep and pathological features. While both MILs offer patch-level interpretability, only ours provides pathological feature-level rationale for WSI predictions. The attention maps in SI-MIL are grounded on geometrically and physically-interpretable descriptors.
In computational pathology, efforts to achieve WSI-level interpretability have predominantly focused on two directions: (1) identifying salient regions in a WSI, and (2) employing post-hoc techniques to elucidate the underlying patterns in salient regions. The first approach, employed by traditional MIL, includes techniques such as visualization of attention maps [21, 36, 56, 68, 57, 67] and post-hoc gradient-based saliency [70, 47], which highlight image patches that influence the model prediction. These techniques, though useful, may not offer a complete understanding of the model’s decisions and can result in visualizations that are hard for experts to interpret due to a lack of user-friendly feature grounding [4]. The latter approach involves extracting interpretable pathological features from the MIL-identified salient patches and then conducting statistical analyses to find correlations between these features and the WSI ground truths in a post-hoc manner [43, 10, 15]. However, there is a clear disconnect between the deep features used for MIL training and the pathological features. Using post-hoc pathological features to explain deep models can be sub-optimal [51]. Furthermore, patches with high attention may be crucial for deep feature space, but may not be optimal in pathological feature space, thus compromising interpretability.
@@
The answer is yes. In this paper, we provide the first WSI solution with both inherent interpretability and strong prediction power. Our key observation is that a highly accurate deep model is not unique; there can be many optimal or close-to-optimal deep models for a dataset/task, due to over-parameterization [45, 27, 30]. Therefore, we hypothesize that we can alter the learning procedure and find an alternative model with desired interpretability and still be powerful in prediction. In particular, we propose to pair a deep MIL model with an interpretable model grounded on pathological features during training. Through co-learning, the MIL retains its predictive power. Meanwhile, it is sufficiently “tamed” by the co-learned interpretable model, which renders interpretability. As shown in Fig. 1, the tamed deep MIL model has a different attention map from the standard MIL. It is attending to patches which can also be discriminated by the companion interpretable model.
Our method, Self-Interpretable MIL (SI-MIL), is a dual branch network, consisting of a conventional MIL and a novel Self-Interpretable (SI) branch. The MIL exploits deep features’ discriminative power to guide the SI branch. Grounded on pathological features, the SI branch then provides a linear prediction. A differentiable Top-$K$ operator for selecting patches, connects the two branches and enables end-to-end co-learning. To highlight, SI-MIL is inherently interpretable [51, 5] due to the linear mapping between the pathological features and the output predictions. Therefore, it can reflect the impact of each feature on the output, thus providing a feature-level rationale, as shown in Fig. 1. Also, by leveraging the potential of a deep feature extractor, state-of-the-art MIL, and geometrically and physically- interpretable pathological features, SI-MIL counters the well-known myth of unavoidable model interpretability and performance trade-off [4, 51]. Notably, SI-MIL is generic enough to substitute any state-of-the-art MIL method in the MIL branch. In summary, our main contributions are:
@@
  • We provide a comprehensive dataset for $\sim$2.2K WSIs, featuring nuclei maps, pathological features, and SI-MIL derived outputs, with the aim of streamlining the resource intensive preprocessing towards interpretability studies in computational pathology.
@@ 2 Related work
Post-hoc interpretability methods: These methods fall into two categories: patch-level and WSI-level. Patch-level techniques, like GradCAM and Layer-Wise Relevance Propagation (used in [19, 53]), highlight key pixels in model predictions. For deeper insight, studies like [22, 23] use biological entity-based graphs for pathologist-friendly explanations. At the WSI-level, interpretability is primarily achieved through attention maps that identify salient regions in WSIs. Additionally, few methods, such as those by  [70, 47], use segmentation maps or gradient-based techniques to localize significant areas. However, these methods, as [51] notes, may suffer from a disconnect from the model’s computations. In pathology, this is particularly evident when comparing the deep features used for MIL training and the pathological features used for subsequent analysis  [43, 10, 15], revealing a disparity in the features for training versus those for feature correlation.
@@ 3 Method
Figure 2: Overview of SI-MIL: Conventional MIL branch guides the Patch Attention-Guided Top-$K$ (PAG Top-K) patch selection module to select the pathological features of key regions from WSI, followed by linear scaling in the Self-Interpretable branch, and linear prediction.
In this section, we present the details of our dual branch SI-MIL (overview in Fig. 2), consisting of a conventional MIL branch and a Self-Interpretable (SI) branch, for analyzing WSIs. We describe the conventional MIL in Sec. 3.1 and present the complete SI-MIL framework in Sec. 3.2. Detailed description of the feature extraction pipelines, i.e., the process of extracting black-box deep features ($g$) and interpretable pathological features ($f$), is provided in Sec. 3.3.
@@ 3.1 Conventional MIL
[…]
@@
$\hat{Y}_{g}=\psi\Big{(}\sum_{i=1}^{N}C(\alpha_{i}\cdot\tilde{g}_{i})\Big{)}$ (2)
@@
3.2 Self-Interpretable MIL (SI-MIL)
As shown in Fig. 2a, along with the aforementioned conventional MIL as a branch, SI-MIL consists of a Patch Attention-Guided Top-$K$ (PAG Top-$K$) module and a SI branch. The PAG Top-$K$ module aims for a differentiable selection of top $K$ patches identified in the MIL branch; thus enabling the co-learning with the SI branch. This branch operates on these top $K$ patches, by leveraging a feature attention module to linearly scale the corresponding relevant pathological features. The patch-wise pathological features and feature attention scores are subsequently aggregated by a linear predictor for slide-level task. SI-MIL denotes the dual branch co-learning framework that discriminates complex WSIs using a linear equation, advancing interpretability by introducing feature-level insights while maintaining high performance and complementing existing MILs. The details of the individual components are described in the following sections.
PAG Top-$K$ patch selection module: This module leverages the patch attention scores $\alpha$ from the conventional MIL branch to select the $K$ most salient patches in a WSI. As the naive Top-$K$ operation is non-differentiable, we use the differentiable perturbed Top-$K$ operation from [11, 58]. This perturbed Top-$K$ operation is imperative to enable the co-learning of both the branches: conventional MIL and SI branch. Following patch selection, only the pathological features of the salient patches are utilized in the subsequent steps. Therefore, the use of deep features in the MIL branch does not hinder the interpretability of SI-MIL; rather it guides the selection of informative patches, which is denoted as:
$S_{K}=\text{TopK}(\alpha,K)$ (3)
where $S_{K}$ denotes the indices of the selected top $K$ patches.
Feature Attention module: The $A^{f}(\cdot)$ module consists of a patch feature mixing network and gated attention network. Their synergistic integration forms a learnable feature selector without interfering with the interpretability of SI-MIL. First, the pathological feature matrix $M\in\mathbb{R}^{K\times d}$, corresponding to the $S_{K}$ patches, is transposed and fed to a patch feature mixing network PF-Mixer, ${PF}(\cdot)$. It contextualizes each value in $M^{T}$ with the top $K$ patches and $d$ features. In practice, ${PF}(\cdot)$ is implemented via MLP layers [59], with separate layers dedicated to mixing spatial patch information and per-patch feature information. Subsequently, gated attention network $G(\cdot)$ processes each row of the matrix $M^{T}\in\mathbb{R}^{d\times K}$ independently to determine the attention score $\beta_{j}$ for each feature $d_{j}$, computed as:
$\beta_{j}=G({PF}(M^{T}));\hskip 14.22636ptj\in\{1,2,...d\}$ (4)
To enforce the model to be dependent on most salient features, we scale the feature attention scores $\beta$ as follows: $\beta$ values are first scaled using percentile $Pr_{\gamma}$ (where $\gamma$ is the $\gamma^{th}$ precentile) and standard deviation (std), and then sigmoid activated with a hyper-parameter, temperature ($t$) as shown in Eqn. 5. This operation enforces the $\beta$ values above $Pr_{\gamma}$ towards 1 and remaining towards 0, thereby imposing sparsity in feature selection.
[…]
These feature attention values are used to linearly scale the pathological feature matrix $M$ such that the salient features are emphasized while attenuating others:
$M^{{}^{\prime}}_{ij}=\beta_{j}\times M_{ij};\hskip 14.22636pti\in\{1,2,...K\};\hskip 2.84544ptj\in\{1,2,...d\}$ (6)
Linear Predictor and Aggregation: Following the attention scaling of the pathological features corresponding to the $S_{K}$ patches, the features are fed to a linear predictor $L(\cdot)$ characterized by weights $w(\cdot)$ and bias $b$ as:
$M^{{}^{\prime\prime}}_{i}=\sum_{j=1}^{d}w_{j}M^{{}^{\prime}}_{ij}+b;\hskip 14.22636pti\in\{1,2,...K\}$ (7)
$\hat{Y}_{f}=\psi\Big{(}\sum_{i=1}^{K}M^{{}^{\prime\prime}}_{i}\Big{)}$ (8)
It can be observed that the WSI-level prediction in the SI branch can be decomposed into a linear combination of feature attention scores $\beta$, classifier weights $w(\cdot)$, and the pathological feature matrix $M\in\mathbb{R}^{K\times d}$ of the top $K$ patches ($S_{K}$), given as:
$\hat{Y}_{f}=\psi\Big{(}\sum_{i=1}^{K}\sum_{j=1}^{d}w_{j}\beta_{j}M_{ij}+b\Big{)}$ (9)
Optimization: Given the true label $Y$ for a WSI, the predictions $\hat{Y}_{g}$ from the MIL branch and $\hat{Y}_{f}$ from the SI branch, SI-MIL is optimized using slide-level cross entropy losses $\mathcal{L}_{CE}$ for both the predictions with respect to the true label. This joint optimization tames the patch attention module to select the patches collaboratively in the deep feature and pathological feature space. To enhance the performance of the SI branch, a knowledge distillation loss $\mathcal{L}_{KD}$ is optimized based on the mean squared error between $\hat{Y}_{f}$ and $\hat{Y}_{g}$. $\mathcal{L}_{KD}$ enforces alignment in performance between the two branches. The overall loss is computed as:
Note that during inference, prediction from any branch can be used. However to enforce interpretability, the WSI-level prediction is obtained from the SI branch, i.e., $\hat{Y}_{f}$ is considered for slide-level prediction and the non-interpretable branch’s output $\hat{Y}_{g}$ is discarded.
@@ 3.3 WSI patch feature extraction
    Hand-crafted pathological features: We use HoVer-Net [17], pretrained on PanNuke [16] dataset, to segment and classify nuclei into 5 classes in each $p_{i}$. Then, pathologist-friendly features $f_{i}\in\mathbb{R}^{d}$ are extracted to quantify nuclei morphology and spatial distribution properties in $p_{i}$. These features can be grouped as:
@@ 4.1 Datasets and Implementation details
Patch and feature extraction: Patches of size $224\times 224$ at 5$\times$ magnification and corresponding $1792\times 1792$ at 40$\times$ are extracted for each dataset. For deep features extraction, we pretrain ViT-S [13] with DINO [8] on the 5$\times$ patches from the training splits of individual datasets mentioned above. Pathological features are extracted on corresponding patches at 40×40$\times$40 ×.
MIL setting: Additive ABMIL [24] is adopted as the conventional MIL in this study. $A^{p}(\cdot)$ and $A^{f}(\cdot)$ are gated attention modules adopted from [36]. For all the MIL experiments, the batch size is set to 1 to handle WSIs of variable bag sizes. For robustness, 5-fold cross-validation is performed on the train split and the mean performance on the held-out test split is reported. More implementation details are provided in the Supp. Sec. 8. Note that SI-MIL is evaluated with only DINO ViT-S features. Experimentation with other deep features is left for future exploration.

### v2

@@ Abstract
Introducing interpretability and reasoning into Multiple Instance Learning (MIL) methods for Whole Slide Image (WSI) analysis is challenging, given the complexity of gigapixel slides. Traditionally, MIL interpretability is limited to identifying salient regions deemed pertinent for downstream tasks, offering little insight to the end-user (pathologist) regarding the rationale behind these selections. To address this, we propose Self-Interpretable MIL (SI-MIL), a method intrinsically designed for interpretability from the very outset. SI-MIL employs a deep MIL framework to guide an interpretable branch grounded on handcrafted pathological features, facilitating linear predictions. Beyond identifying salient regions, SI-MIL uniquely provides feature-level interpretations rooted in pathological insights for WSIs. Notably, SI-MIL, with its linear prediction constraints, challenges the prevalent myth of an inevitable trade-off between model interpretability and performance, demonstrating competitive results compared to state-of-the-art methods on WSI-level prediction tasks across three cancer types. In addition, we thoroughly benchmark the local- and global-interpretability of SI-MIL in terms of statistical analysis, a domain expert study, and desiderata of interpretability, namely, user-friendliness and faithfulness. Code and Dataset is available at: github.com/bmi-imaginelab/SI-MIL
@@ 1 Introduction
Figure 1: Unlike conventional MIL, SI-MIL co-learns from deep and handcrafted features (referred to as PathExpert features). While both MILs offer patch-level interpretability, only ours provides PathExpert feature-level rationale for WSI predictions. The attention maps in SI-MIL are grounded on geometrically and physically-interpretable descriptors.
In computational pathology, efforts to achieve WSI-level interpretability have predominantly focused on two directions: (1) identifying salient regions in a WSI, and (2) employing post-hoc techniques to elucidate the underlying patterns in salient regions. The first approach, employed by traditional MIL, includes techniques such as visualization of attention maps [21, 36, 56, 68, 57, 67] and post-hoc gradient-based saliency [70, 47], which highlight image patches that influence the model prediction. These techniques, though useful, may not offer a complete understanding of the model’s decisions and can result in visualizations that are hard for experts to interpret due to a lack of user-friendly feature grounding [4]. The latter approach involves extracting interpretable handcrafted features (henceforth referred to as PathExpert features) from the MIL-identified salient patches and then conducting statistical analyses to find correlations between these features and the WSI ground truths in a post-hoc manner [43, 10, 15]. However, there is a clear disconnect between the deep features used for MIL training and the PathExpert features. Using post-hoc PathExpert features to explain deep models can be sub-optimal [51]. Furthermore, patches with high attention may be crucial for deep feature space, but may not be optimal in PathExpert feature space, thus compromising interpretability.
@@
The answer is yes. In this paper, we provide the first WSI solution with both inherent interpretability and strong prediction power. Our key observation is that a highly accurate deep model is not unique; there can be many optimal or close-to-optimal deep models for a dataset/task, due to over-parameterization [45, 27, 30]. Therefore, we hypothesize that we can alter the learning procedure and find an alternative model with desired interpretability and still be powerful in prediction. In particular, we propose to pair a deep MIL model with an interpretable model grounded on PathExpert features during training. Through co-learning, the MIL retains its predictive power. Meanwhile, it is sufficiently “tamed” by the co-learned interpretable model, which renders interpretability. As shown in Fig. 1, the tamed deep MIL model has a different attention map from the standard MIL. It is attending to patches which can also be discriminated by the companion interpretable model.
Our method, Self-Interpretable MIL (SI-MIL), is a dual branch network, consisting of a conventional MIL and a novel Self-Interpretable (SI) branch. The MIL exploits deep features’ discriminative power to guide the SI branch. Grounded on PathExpert features, the SI branch then provides a linear prediction. A differentiable Top-$K$ operator for selecting patches, connects the two branches and enables end-to-end co-learning. To highlight, SI-MIL is inherently interpretable [51, 5] due to the linear mapping between the PathExpert features and the output predictions. Therefore, it can reflect the impact of each feature on the output, thus providing a feature-level rationale, as shown in Fig. 1. Also, by leveraging the potential of a deep feature extractor, state-of-the-art MIL, and geometrically and physically- interpretable PathExpert features, SI-MIL counters the well-known myth of unavoidable model interpretability and performance trade-off [4, 51]. Notably, SI-MIL is generic enough to substitute any state-of-the-art MIL method in the MIL branch. In summary, our main contributions are:
@@
  • We provide a comprehensive dataset for $\sim$2.2K WSIs, featuring nuclei maps, PathExpert features, and SI-MIL derived outputs, with the aim of streamlining the resource intensive preprocessing towards interpretability studies in computational pathology.
@@ 2 Related work
Post-hoc interpretability methods: These methods fall into two categories: patch-level and WSI-level. Patch-level techniques, like GradCAM and Layer-Wise Relevance Propagation (used in [19, 53]), highlight key pixels in model predictions. For deeper insight, studies like [22, 23] use biological entity-based graphs for pathologist-friendly explanations. At the WSI-level, interpretability is primarily achieved through attention maps that identify salient regions in WSIs. Additionally, few methods, such as those by  [70, 47], use segmentation maps or gradient-based techniques to localize significant areas. However, these methods, as [51] notes, may suffer from a disconnect from the model’s computations. In pathology, this is particularly evident when comparing the deep features used for MIL training and the handcrafted features used for subsequent analysis  [43, 10, 15], revealing a disparity in the features for training versus those for feature correlation.
@@ 3 Method
Figure 2: Overview of SI-MIL: Conventional MIL branch guides the Patch Attention-Guided Top-$K$ (PAG Top-K) patch selection module to select the PathExpert features of key regions from WSI, followed by linear scaling in the Self-Interpretable branch, and linear prediction.
In this section, we present the details of our dual branch SI-MIL (overview in Fig. 2), consisting of a conventional MIL branch and a Self-Interpretable (SI) branch, for analyzing WSIs. We describe the conventional MIL in Sec. 3.1. Detailed description of the feature extraction pipelines, i.e., the process of extracting black-box deep features ($g$) and interpretable PathExpert features ($f$), is provided in Sec. 3.2. Finally, we present the complete SI-MIL framework in Sec. 3.3.
@@ 3.1 Conventional MIL
$\vspace{-0.2cm}\tilde{g}_{i}=H(g_{i});\hskip 14.22636pt\alpha_{i}=A^{p}(\tilde{g}_{i});\hskip 14.22636pti\in\{1,2,...N\}$ (1)
@@
$\vspace{-0.2cm}\hat{Y}_{g}=\psi\Big{(}\sum_{i=1}^{N}C(\alpha_{i}\cdot\tilde{g}_{i})\Big{)}$ (2)
@@ 3.2 WSI patch feature extraction
    Hand-crafted PathExpert features: We use HoVer-Net [17], pretrained on PanNuke [16] dataset, to segment and classify nuclei into 5 classes in each $p_{i}$. Then, pathologist-friendly features $f_{i}\in\mathbb{R}^{d}$ are extracted to quantify nuclei morphology and spatial distribution properties in $p_{i}$. These features can be grouped as:
@@
3.3 Self-Interpretable MIL (SI-MIL)
As shown in Fig. 2a, along with the aforementioned conventional MIL as a branch, SI-MIL consists of a Patch Attention-Guided Top-$K$ (PAG Top-$K$) module and a SI branch. The PAG Top-$K$ module aims for a differentiable selection of top $K$ patches identified in the MIL branch; thus enabling the co-learning with the SI branch. This branch operates on these top $K$ patches, by leveraging a feature attention module to linearly scale the corresponding relevant PathExpert features. The patch-wise PathExpert features and feature attention scores are subsequently aggregated by a linear predictor for slide-level task. SI-MIL denotes the dual branch co-learning framework that discriminates complex WSIs using a linear equation, advancing interpretability by introducing feature-level insights while maintaining high performance and complementing existing MILs. The details of the individual components are described in the following sections.
PAG Top-$K$ patch selection module: This module leverages the patch attention scores $\alpha$ from the conventional MIL branch to select the $K$ most salient patches in a WSI. As the naive Top-$K$ operation is non-differentiable, we use the differentiable perturbed Top-$K$ operation from [11, 58]. This perturbed Top-$K$ operation is imperative to enable the co-learning of both the branches: conventional MIL and SI branch. Following patch selection, only the PathExpert features of the salient patches are utilized in the subsequent steps. Therefore, the use of deep features in the MIL branch does not hinder the interpretability of SI-MIL; rather it guides the selection of informative patches, which is denoted as:
$\vspace{-0.2cm}S_{K}=\text{TopK}(\alpha,K)$ (3)
where $S_{K}$ denotes the indices of the selected top $K$ patches.
Feature Attention module: The $A^{f}(\cdot)$ module consists of a patch feature mixing network and gated attention network. Their synergistic integration forms a learnable feature selector without interfering with the interpretability of SI-MIL. First, the PathExpert feature matrix $M\in\mathbb{R}^{K\times d}$, corresponding to the $S_{K}$ patches, is transposed and fed to a patch feature mixing network PF-Mixer, ${PF}(\cdot)$. It contextualizes each value in $M^{T}$ with the top $K$ patches and $d$ features. In practice, ${PF}(\cdot)$ is implemented via MLP layers [59], with separate layers dedicated to mixing spatial patch information and per-patch feature information. Subsequently, gated attention network $G(\cdot)$ processes each row of the matrix $M^{T}\in\mathbb{R}^{d\times K}$ independently to determine the attention score $\beta_{j}$ for each feature $d_{j}$, computed as:
$\vspace{-0.2cm}\beta_{j}=G({PF}(M^{T}));\hskip 14.22636ptj\in\{1,2,...d\}$ (4)
To enforce the model to be dependent on most salient features, we scale the feature attention scores $\beta$ as follows: $\beta$ values are first scaled using percentile $Pr_{\gamma}$ (where $\gamma$ is the $\gamma^{th}$ precentile) and standard deviation (std), and then sigmoid activated with a hyper-parameter, temperature ($t$) as shown in Eqn. 5. This operation enforces the $\beta$ values above $Pr_{\gamma}$ towards 1 and remaining towards 0, thereby imposing sparsity in feature selection. Note that for brevity, we denote the scaled values of $\beta$ with same notation in Eqn. 5.
$\vspace{-0.2cm}\beta_{j}=\frac{\beta_{j}-Pr_{\gamma}(\beta)}{\text{std}(\beta)};\hskip 14.22636pt\beta_{j}=\frac{1}{1+e^{-\beta_{j}\times t}}$ (5)
These feature attention values are used to linearly scale the PathExpert feature matrix $M$ such that the salient features are emphasized while attenuating others:
$\vspace{-0.2cm}M^{{}^{\prime}}_{ij}=\beta_{j}\times M_{ij};\hskip 14.22636pti\in\{1,2,...K\};\hskip 2.84544ptj\in\{1,2,...d\}$ (6)
Linear Predictor and Aggregation: Following the attention scaling of the PathExpert features corresponding to the $S_{K}$ patches, the features are fed to a linear predictor $L(\cdot)$ characterized by weights $w(\cdot)$ and bias $b$ as:
$\vspace{-0.2cm}M^{{}^{\prime\prime}}_{i}=\sum_{j=1}^{d}w_{j}M^{{}^{\prime}}_{ij}+b;\hskip 14.22636pti\in\{1,2,...K\}$ (7)
$\vspace{-0.1cm}\hat{Y}_{f}=\psi\Big{(}\sum_{i=1}^{K}M^{{}^{\prime\prime}}_{i}\Big{)}$ (8)
It can be observed that the WSI-level prediction in the SI branch can be decomposed into a linear combination of feature attention scores $\beta$, classifier weights $w(\cdot)$, and the PathExpert feature matrix $M\in\mathbb{R}^{K\times d}$ of the top $K$ patches ($S_{K}$), given as:
$\vspace{-0.2cm}\hat{Y}_{f}=\psi\Big{(}\sum_{i=1}^{K}\sum_{j=1}^{d}w_{j}\beta_{j}M_{ij}+b\Big{)}$ (9)
Optimization: Given the true label $Y$ for a WSI, the predictions $\hat{Y}_{g}$ from the MIL branch and $\hat{Y}_{f}$ from the SI branch, SI-MIL is optimized using slide-level cross entropy losses $\mathcal{L}_{CE}$ for both the predictions with respect to the true label. This joint optimization tames the patch attention module to select the patches collaboratively in the deep feature and PathExpert feature space. To enhance the performance of the SI branch, a knowledge distillation loss $\mathcal{L}_{KD}$ is optimized based on the mean squared error between $\hat{Y}_{f}$ and $\hat{Y}_{g}$. $\mathcal{L}_{KD}$ enforces alignment in performance between the two branches. The overall loss is computed as:
where $\lambda$ is used as a weight to align the scale of $\mathcal{L}_{KD}$ with the $\mathcal{L}_{CE}$ losses of deep feature and self-interpretable branch. Note that during inference, prediction from any branch can be used. However to enforce interpretability, the WSI-level prediction is obtained from the SI branch, i.e., $\hat{Y}_{f}$ is considered for slide-level prediction and the non-interpretable branch’s output $\hat{Y}_{g}$ is discarded.
@@ 4.1 Datasets and Implementation details
Patch and feature extraction: Patches of size $224\times 224$ at 5$\times$ magnification and corresponding $1792\times 1792$ at 40$\times$ are extracted for each dataset. For deep features extraction, we pretrain ViT-S [13] with DINO [8] on the 5$\times$ patches from the training splits of individual datasets mentioned above. PathExpert features are extracted on corresponding patches at 40×40$\times$40 ×.
MIL setting: Additive ABMIL [24] is adopted as the conventional MIL in this study. $A^{p}(\cdot)$ and $A^{f}(\cdot)$ are deep neural network based gated attention modules adopted from [36]. For all the MIL experiments, the batch size is set to 1 to handle WSIs of variable bag sizes. For robustness, 5-fold cross-validation is performed on the train split and the mean performance on the held-out test split is reported. By default, ##\##PF-Mixer layers $=$4, $\lambda=20$, $K=20$, $\gamma=0.75$, and $t=3$. More implementation details are provided in the Supp. Sec. 9. Note that SI-MIL is evaluated with only DINO ViT-S features. Experimentation with other deep features is left for future exploration.